# Conformalized Generative Bayesian Imaging: An Uncertainty Quantification Framework for Computational Imaging

### v1

@@ Abstract
Index Terms:
Uncertainty quantification, generative models, Bayesian neural networks, conformal prediction, image reconstruction, inverse problems, epistemic uncertainty, aleatoric uncertainty, posterior sampling.
@@ II-A Problem Setup
For such inverse problems, throughout the remainder of this paper, we make two main assumptions about the problem setup. First, we assume access to a training dataset $\mathcal{D}$ containing measurement vectors and corresponding reference images. Second, we assume that we already have a generative model-based posterior sampling method at hand that aims to generate samples from the posterior distribution of the underlying image given measurements $p_{\boldsymbol{\mathsf{x}}|\boldsymbol{\mathsf{y}}}(\boldsymbol{\mathsf{x}}|\mathbf{y})$. In the subsequent sections, we denote this generative model-based posterior sampling method by the tuple $(G,p_{\boldsymbol{\mathsf{z}}},\mathcal{T})$. In this notation, $G:\mathbb{C}^{M}\times\mathbb{R}^{Z}\to\mathbb{C}^{N}$ is a conditional generative model generating the samples; $\boldsymbol{\mathsf{z}}\sim p_{\boldsymbol{\mathsf{z}}}$ is an $\mathbb{R}^{Z}$-valued random latent variable; and $\mathcal{T}$ is the training procedure followed by the generative model-based posterior sampling method to train the generative model $G$. These assumptions are often justifiable for a variety of imaging problems, although they may not be plausible for certain imaging problems, especially if obtaining reference images is unfeasible.
@@ II-B Proposed Uncertainty Quantification Approach
[…]
@@
where $(\cdot)^{\top}$ denotes the Hermitian transpose operator. By following the uncertainty decomposition idea presented in [53], the proposed framework can decompose the predictive uncertainty estimate in (7) into epistemic and aleatoric uncertainty estimates as follows:
@@ III-A1 CT Experiments
For the computed tomography (CT) experiments, we obtained $11420$ $512\times 512$ reference images from the LUNA dataset [63] and resized each reference image to $256\times 256$ pixels. Then, we normalized each reference image such that the interval $[-1000,3000]$ Hounsfield unit (HU) was mapped into the interval $[0,1]$. We used $11220$ of those references images for the training dataset and split the remaining ones into two parts to be used for the validation and test datasets, each containing $100$ reference images. For each reference image in the training, validation, and test datasets, we generated the corresponding measurements by calculating its Radon transform with $72$ views (corresponding to approximately 5×5$\times$5 × dose reduction) and adding white Gaussian noise such that signal-to-noise ratio was approximately $50$ decibels.
We used a generative adversarial network-based posterior sampling method called deep posterior sampling [22] (DPS) to build the proposed framework. At the training stage of the proposed framework, we trained $T_{2}=5$ copies (initialized with different random weights) of the conditional Wasserstein generative adversarial network proposed by the DPS method for $10$ epochs by following the training procedure described in [22], except that we set the mini-batch size to $16$ and did not use any learning rate decay. We used the validation dataset to tune the hyperparameters and monitor the individual performance of each DPS instance in the ensemble. At the inference stage, for a given test measurement vector, we first calculated the filtered backprojection of the test measurement vector and then used it as an input to all generative models in the ensemble together with a sample from the prior distribution of the latent variable. We repeated this process $T_{1}=128$ times and obtained the corresponding reconstructed images and uncertainty maps.
@@ III-A2 MRI Experiments
For the magnetic resonance imaging (MRI) experiments, we obtained $70372$ $256\times 256$ reference images from the IXI dataset [64] and normalized each image such that the intensity values of each image lie in the interval $[0,1]$. We used $70172$ of the reference images for the training dataset and split the remaining reference images into two sets to be used for the validation and test datasets. Both the validation and test datasets contain $100$ reference images. For each reference image in the training, validation, and test datasets, we generated the measurements by calculating its subsampled Fourier transform and adding white Gaussian noise. We used the radial subsampling pattern and collected only $20\%$ of the Fourier coefficients, corresponding to approximately 5×5$\times$5 × acceleration. We fixed the standard deviation of the additive Gaussian noise such that the SNR was approximately $50$ decibels. It is worth mentioning that we employed the IXI dataset due to its convenience for experimentation. While it may be processed or oversimplified for real-world MR imaging, our goal is not to demonstrate state-of-the-art MRI reconstruction performance on real-world data, but rather to evaluate the relative improvements offered by the proposed framework in comparison to the generative model-based posterior sampling methods used within the proposed framework.
We built the proposed method on a conditional variational autoencoder (CVAE) [25]-based posterior sampling method. We utilized the U-Net [65] architecture used in [66] to obtain the baseline network of the CVAE and used its dual-head version to create the prior network, generation network, and the recognition network (see [25] for details on the roles of these networks) of the CVAE. At the training stage of the proposed framework, we trained $T_{2}=5$ instances of the CVAE-based posterior sampling method, initialized with different random weights. We trained each instance for $5$ epochs by minimizing the empirical lower bound provided in [25, Eq. (5)]. We set the mini-batch size to $16$ and used the Adam [67] optimizer with the learning rate of $10^{-4}$. At the inference stage, for a given test measurement vector, we first performed zero-filling and then used the result as an input to each CVAE in the ensemble together with a sample from the prior network. We repeated this procedure $T_{1}=128$ times and obtained the corresponding reconstructed images and uncertainty maps.
Figure 1: Variation in the epistemic uncertainty estimates offered by the proposed framework for a randomly chosen CT test measurement vector as the number of examples in the training dataset is changed. For this example, the generative model-based posterior sampling method employed by the proposed framework is deep posterior sampling [22] (see III-A for further details).
@@ III-A3 Image Inpainting Experiments
For the image inpainting experiments presented in Section III-F, we obtained $60025$ $32\times 32$ reference images from the MNIST dataset [68] and normalized them linearly such that their entries lie in the interval $[-1,1]$. We then split the reference images into two sets to be used for the training and test datasets, having $60000$ and $25$ samples, respectively. For each reference image in the test dataset, we generated its corresponding measurement vector by multiplying the image with a mask and adding white Gaussian noise. We used a mask that randomly samples $10\%$ of the image pixels and fixed the standard deviation of the noise to $0.05$.
We built the proposed framework on top of a diffusion model-based posterior sampling method called diffusion posterior sampling [14], whose open-source implementation is provided in [69]. We trained the diffusion model [70] used within the diffusion posterior sampling technique for $4\times 10^{4}$ iterations by modifying the open-source implementation provided in [71]. At the inference stage, we used each test measurement vector as an input to the diffusion posterior sampling together with $T_{1}=32$ samples from the latent distribution. We repeated this procedure $T_{2}=5$ times and obtained the corresponding reconstructions and the uncertainty maps.
@@ III-B Impact of the Training Dataset Size on Epistemic Uncertainty Estimates
Figure 3: Effect of the test measurement vectors that are not well-represented by the training dataset on the epistemic uncertainty estimates provided by the proposed framework. The first row shows the output of filtered backprojection (FBP), the output of zero-filling (ZF), the predictive mean provided by the proposed framework, and the epistemic uncertainty estimate offered by the proposed framework for both the CT and MRI problems, where there are no abnormalities present in the test measurement vectors. The second row shows the results of the cases where abnormalities are introduced into the same test measurement vectors (please refer to Section III-C for the details of these abnormalities).
@@ III-C Epistemic Uncertainty Estimates and Abnormalities Occurring at the Inference Time
For the CT problem, we inserted two synthetic metal implants on a test reference image by following the simulation procedure described in [72, 73] and generated the corresponding test measurement vector by following the procedure described in Section III-A. We then used the resulting test measurement vector as an input to the proposed framework and obtained the corresponding reconstructed image and epistemic uncertainty map. Similarly, for the MRI problem, we introduced an abnormality to a test measurement vector by adding random spikes on the Fourier transform coefficients, which is sometimes referred to as the Herringbone artifact, using the TorchIO toolbox [74]. We then used the final test measurement vector as an input to the proposed framework.
Figure 3 displays the results for both the CT and MRI problems, as well as the results for the reference cases where no abnormalities are present on the test measurement vectors. By comparing the first and the fourth columns of Figure 3, we see that the introduced abnormalities caused visually apparent deviations on the outputs of the FBP and ZF methods. Since the outputs of the FBP and ZF methods are essentially what the generative models used within the WGAN-based proposed framework instance and the CVAE-based proposed framework instance are conditioned upon, by examining the second and the fifth columns of Figure 3, we observe that the reconstructed images provided by the proposed framework contain artifacts for the cases where the test measurement vectors are not well-represented by the CT and MRI training datasets. However, by looking at the epistemic uncertainty estimates shown in the third and sixth columns of Figure 3, we see that the proposed framework has clearly identified the abnormalities on the test measurement vector as well as the artifacts caused by those abnormalities. These two examples highlight that although the proposed framework is not capable of successfully recovering the underlying image from a test measurement that is not well-represented by the training dataset, it offers a mechanism to identify and detect such problematic cases.
@@ III-D Quality of the Predictive Uncertainty Estimates
In this section, we assess the quality of the final uncertainty estimates offered by the proposed framework, which integrates both aleatoric and epistemic uncertainties, and compare it to the final uncertainty estimates produced by the generative model-based posterior sampling method used within the proposed framework, which quantifies only the aleatoric uncertainty. For the CT problem, we compare the proposed framework with the original DPS method. Similarly, we compare the proposed framework with the original CVAE-based posterior sampling method for the MRI problem. Because we have $T_{2}=5$ different instances of the original DPS method and the CVAE-based posterior sampling method at hand, we also compare each instance to one another to examine the variations on the quality of the uncertainty estimates induced by using different parameter values for the generative model-based posterior sampling methods.
In our experiments, we assessed the quality of the final uncertainty estimates through the negative predictive log-likelihood metric (details provided in the supplementary material). Figure 4 presents the negative predictive log-likelihood (NPLL) values of the evaluated methods for the CT and MRI problems. Careful examination of the figure provides several key observations about the predictive performance of the evaluated methods. First, it highlights that the quality of the predictive uncertainty estimates of both the DPS method and the CVAE-based posterior sampling method shows notable variations depending on the initializations of the parameters. Secondly, we see that the proposed framework achieves the lowest NPLL values across all evaluated methods, demonstrating superior predictive performance compared to the DPS method and the CVAE-based posterior sampling method. These two observations demonstrate the advantage of utilizing multiple realizations of the parameters during the inference stage, as opposed to methods that rely on a single realization of their parameters.
Figure 4: Negative predictive log-likelihood (NPLL) results for deep posterior sampling [22] (DPS), the conditional variational autoencoder [25]-based posterior sampling method (CVAE), and the proposed deep ensembling [41]-based framework (DE). The results are presented for the CT (top) and MRI (bottom) problems.
@@ III-E Reconstruction Performance
The objective of this subsection is to assess the quality of the reconstructed images provided by the proposed framework and compare it against relevant baseline methods. For this purpose, we used structural similarity index [75] (SSIM) as our evaluation metric. When an evaluated method produced a collection of reconstructions for a given measurement vector rather than a single reconstructed image, we calculated the SSIM between the mean of the collection and the reference image.
For the CT problem, we compared the performance of the proposed framework against filtered backprojection (FBP), a state-of-the-art deep learning-based image reconstruction method FBPConvNet [66], and the instances of the DPS method used within the proposed framework. For the MRI problem, we compared the performance of the proposed framework against zero-filling (ZF), a state-of-the-art deep learning-based image reconstruction method called BPConvNet [66], and the instances of the CVAE-based posterior sampling method used within the proposed framework. Figure 5 displays the resulting values of each metric calculated across the corresponding test dataset for different reconstruction methods. The visual results can be found in the supplementary material.
@@
Figure 5: SSIM results for filtered backprojection (FBP), FBPConvNet [66] (FBPCN), deep posterior sampling [22] (DPS), zero-filling (ZF), BPConvNet [66] (BPCN), the conditional variational autoencoder [25]-based posterior sampling method (CVAE), and the proposed deep ensembling [41]-based framework (DE). Results shown for the CT (top) and MRI (bottom) problems.
@@ III-F Computationally Efficient Ensembling
Figure 6: SSIM and negative predictive log-likelihood (NPLL) results for diffusion posterior sampling [14] (DiPS), proposed framework with MC Dropout ensembling (MC-DC and MC-DE), deep ensembling (DE), and SWAG ensembling (SE). The results are presented for the image inpainting problem described in Section III-A.

### v2

@@ II-A Problem Setup
For such inverse problems, throughout the remainder of this paper, we make two main assumptions about the problem setup. First, we assume access to a training dataset $\mathcal{D}$ containing measurement vectors and corresponding reference images. Second, we assume that we already have a generative model-based posterior sampling method at hand that is capable of generating samples from the posterior distribution of the underlying image given measurements $p_{\boldsymbol{\mathsf{x}}|\boldsymbol{\mathsf{y}}}(\boldsymbol{\mathsf{x}}|\mathbf{y})$. In the subsequent sections, we denote this generative model-based posterior sampling method by the tuple $(G,p_{\boldsymbol{\mathsf{z}}},\mathcal{T})$. In this notation, $G:\mathbb{C}^{M}\times\mathbb{R}^{Z}\to\mathbb{C}^{N}$ is a conditional generative model generating the samples; $\boldsymbol{\mathsf{z}}\sim p_{\boldsymbol{\mathsf{z}}}$ is an $\mathbb{R}^{Z}$-valued random latent variable; and $\mathcal{T}$ is the training procedure followed by the generative model-based posterior sampling method to train the generative model $G$. These assumptions are often justifiable for a variety of imaging problems, although they may not be plausible for certain imaging problems, especially if obtaining reference images is unfeasible.
@@ II-B Proposed Uncertainty Quantification Approach
$\int_{\mathbb{R}^{P}}\int_{\mathbb{R}^{Z}}p_{\boldsymbol{\mathsf{x}}|\boldsymbol{\mathsf{y}},\boldsymbol{\mathsf{z}},\boldsymbol{\mathsf{\Theta}}}(\boldsymbol{\mathsf{{x}_{*}}}|{\mathbf{y}}_{*},\mathbf{z},\boldsymbol{\theta})p_{\boldsymbol{\mathsf{\Theta}}|\mathsf{D}}(\boldsymbol{\theta}|\mathcal{D})p_{\boldsymbol{\mathsf{z}}}(\mathbf{z})d\mathbf{z}d\boldsymbol{\theta},$ (2)
@@
where $(\cdot)^{\top}$ denotes the transpose operator. By following the uncertainty decomposition idea presented in [53], the proposed framework can decompose the predictive uncertainty estimate in (7) into epistemic and aleatoric uncertainty estimates as follows:
@@ III-A1 CT Experiments
For the computed tomography (CT) experiments, we obtained $11940$ $512\times 512$ reference images from the LUNA dataset [63] and resized each reference image to $256\times 256$ pixels. Then, we normalized each reference image such that the interval [1000,3000][-1000,3000] Hounsfield unit (HU) was mapped into the interval $[0,1]$. We used $11220$ of those reference images for the training dataset and split the remaining ones into two parts to be used for the validation and test datasets, each containing $100$ and $620$ reference images, respectively. For each reference image in the training, validation, and test datasets, we generated the corresponding measurements by calculating its Radon transform with $72$ views (corresponding to approximately $5\times$ dose reduction) and adding white Gaussian noise such that signal-to-noise ratio was approximately $50$ decibels.
We used a generative adversarial network-based posterior sampling method called deep posterior sampling [22] (DPS) to build the proposed framework. To ensure reliable posterior sampling, DPS introduces a novel discriminator formulation that addresses the well-known mode collapse problem commonly observed in generative adversarial networks. At the training stage of the proposed framework, we trained $T_{2}=5$ copies (initialized with different random weights) of the conditional Wasserstein generative adversarial network proposed by the DPS method. We used the validation dataset to tune the hyperparameters and monitor the individual performance of each DPS instance in the ensemble. At the inference stage, for a given test measurement vector, we first calculated the filtered backprojection of the test measurement vector and then used it as an input to all generative models in the ensemble together with a sample from the prior distribution of the latent variable. We repeated this process $T_{1}=128$ times and obtained the corresponding reconstructed images and uncertainty maps.
@@ III-A2 MRI Experiments
For the magnetic resonance imaging (MRI) experiments, we obtained $41877$ $320\times 320$ complex knee MR reference images from the fastMRI dataset [64, 65] and normalized each image such that the intensity values of each magnitude image lie in the interval $[0,1]$. We used $34742$ of the reference images for the training dataset and split the remaining reference images into two sets to be used for the validation and test datasets. The validation and test datasets contain $3521$ and $3614$ reference images, respectively. For each reference image in the training, validation, and test datasets, we generated the measurements by undersampling the full k-space data from the fastMRI dataset. We drew an independent random undersampling mask for each reference image, retaining only $20\%$ of the full k-space coefficients and achieving approximately $5\times$ acceleration.
We built the proposed method on a variational autoencoder [24]-based posterior sampling method proposed in [66], which we refer to as the Uncertainty Quantifying Variational Autoencoder for MRI (UQVAE). To ensure reliable posterior sampling, UQVAE incorporates skip connections in the decoder [67] to mitigate the latent variable collapse problem commonly observed in variational autoencoders. At the training stage of the proposed framework, we trained $T_{2}=5$ instances of the UQVAE method, initialized with different random weights. At the inference stage, for a given test measurement vector, we first performed zero-filling and then used the result as an input to each UQVAE in the ensemble. We repeated this procedure $T_{1}=128$ times and obtained the corresponding reconstructed images and uncertainty maps.
Figure 1: Variation in the epistemic uncertainty estimates offered by the proposed framework for a randomly chosen CT test measurement vector as the number of examples in the training dataset is changed. The fractions of the training data used (0.125, 0.25, 0.5, and 1.0) are indicated above each epistemic uncertainty map. For this example, the generative model-based posterior sampling method employed by the proposed framework is deep posterior sampling [22] (see III-A for further details).
@@ III-A3 Image Inpainting Experiments
For the image inpainting experiments presented in Section III-F, we obtained $60025$ $32\times 32$ reference images from the MNIST dataset [68] and normalized them linearly such that their entries lie in the interval [1,1][-1,1]. We then split the reference images into two sets to be used for the training and test datasets, having $60000$ and $25$ samples, respectively. For each reference image in the test dataset, we generated its corresponding measurement vector by multiplying the image with a mask and adding white Gaussian noise. We used a mask that randomly samples $10\%$ of the image pixels and fixed the standard deviation of the noise to $0.05$.
We built the proposed framework on top of a diffusion model-based posterior sampling method called diffusion posterior sampling [14], whose open-source implementation is provided in [69]. We trained the diffusion model [70] used within the diffusion posterior sampling technique by modifying the open-source implementation provided in [71]. At the inference stage, we used each test measurement vector as an input to the diffusion posterior sampling together with $T_{1}=32$ samples from the latent distribution. We repeated this procedure $T_{2}=5$ times and obtained the corresponding reconstructions and the uncertainty maps. Further implementation details of all methods used in the experiments are provided in the supplementary material to ensure clarity and reproducibility.
@@ III-B Impact of the Training Dataset Size on Epistemic Uncertainty Estimates
Figure 3: Effect of the test measurement vectors that are not well-represented by the training dataset on the epistemic uncertainty estimates provided by the proposed framework. The first row shows the output of filtered backprojection (FBP), the output of zero-filling (ZF), the predictive mean provided by the proposed framework, and the epistemic uncertainty estimate offered by the proposed framework for both the CT and MRI problems, where there are no abnormalities present in the test measurement vectors. The second row shows the results of the cases where abnormalities are introduced into the same test measurement vectors (please refer to Section III-C for the details of these abnormalities). Note: For the MRI problem, epistemic uncertainty maps are computed before the final data consistency layer of UQVAE [66].
@@ III-C Epistemic Uncertainty Estimates and Abnormalities Occurring at the Inference Time
For the CT problem, we inserted two synthetic metal implants on a test reference image by following the simulation procedure described in [72, 73] and generated the corresponding test measurement vector by following the procedure described in Section III-A. We then used the resulting test measurement vector as an input to the proposed framework and obtained the corresponding reconstructed image and epistemic uncertainty map. Similarly, for the MRI problem, we introduced an abnormality to a test measurement vector by adding random spikes on the Fourier transform coefficients, which is sometimes referred to as the Herringbone artifact. We then used the final test measurement vector as an input to the proposed framework.
Figure 3 displays the results for both the CT and MRI problems, as well as the results for the reference cases where no abnormalities are present on the test measurement vectors. By comparing the first and the fourth columns of Figure 3, we see that the introduced abnormalities caused visually apparent deviations on the outputs of the FBP and ZF methods. Since the outputs of the FBP and ZF methods are essentially what the generative models used within the DPS-based proposed framework instance and the UQVAE-based proposed framework instance are conditioned upon, by examining the second and the fifth columns of Figure 3, we observe that the reconstructed images provided by the proposed framework contain artifacts for the cases where the test measurement vectors are not well-represented by the CT and MRI training datasets. However, by looking at the epistemic uncertainty estimates shown in the third and sixth columns of Figure 3, we see that the proposed framework has clearly identified the abnormalities on the test measurement vector as well as the artifacts caused by those abnormalities. These two examples highlight that although the proposed framework is not capable of successfully recovering the underlying image from a test measurement that is not well-represented by the training dataset, it offers a mechanism to identify and detect such problematic cases.
@@ III-D Quality of the Predictive Uncertainty Estimates
In this section, we assess the quality of the final uncertainty estimates offered by the proposed framework, which integrates both aleatoric and epistemic uncertainties, and compare it to the final uncertainty estimates produced by the generative model-based posterior sampling method used within the proposed framework, which quantifies only the aleatoric uncertainty. For the CT problem, we compare the proposed framework with the original DPS method. Similarly, we compare the proposed framework with the original UQVAE method for the MRI problem. Because we have $T_{2}=5$ different instances of the original DPS method and the UQVAE method at hand, we also compare each instance to one another to examine the variations on the quality of the uncertainty estimates induced by using different parameter values for the generative model-based posterior sampling methods.
In our experiments, we assessed the quality of the final uncertainty estimates through the negative predictive log-likelihood metric (details provided in the supplementary material). Figure 4 presents the negative predictive log-likelihood (NPLL) values of the evaluated methods for the CT and MRI problems. Careful examination of the figure provides several key observations about the predictive performance of the evaluated methods. First, it highlights that the quality of the predictive uncertainty estimates of both the DPS method and the UQVAE method shows notable variations depending on the initializations of the parameters. Secondly, we see that the proposed framework achieves the lowest NPLL values across all evaluated methods, demonstrating superior predictive performance compared to the DPS method and the UQVAE method. These two observations demonstrate the advantage of utilizing multiple realizations of the parameters during the inference stage, as opposed to methods that rely on a single realization of their parameters.
Figure 4: Negative predictive log-likelihood (NPLL) results for deep posterior sampling [22] (DPS), the variational autoencoder-based posterior sampling method [66] (UQVAE), and the proposed deep ensembling [41]-based framework (DE). The results are presented for the CT (top) and MRI (bottom) problems.
@@ III-E Reconstruction Performance
The objective of this subsection is to assess the quality of the reconstructed images provided by the proposed framework and compare it against relevant baseline methods. For this purpose, we used structural similarity index [74] (SSIM) as our evaluation metric (results for the mean squared error (MSE) metric are provided in the supplementary material). When an evaluated method produced a collection of reconstructions for a given measurement vector rather than a single reconstructed image, we calculated the SSIM between the mean of the collection and the reference image.
For the CT problem, we compared the performance of the proposed framework against filtered backprojection (FBP), a state-of-the-art deep learning-based image reconstruction method FBPConvNet [75], and the instances of the DPS method used within the proposed framework. For the MRI problem, we compared the performance of the proposed framework against zero-filling (ZF), a state-of-the-art deep learning-based image reconstruction method called BPConvNet [75], and the instances of the UQVAE method used within the proposed framework. Figure 5 displays the resulting values of each metric calculated across the corresponding test dataset for different reconstruction methods. The visual results can be found in the supplementary material.
@@
We also observe that, in the CT experiments, all DPS instances achieve better reconstruction performance than FBPConvNet. We believe that the superior reconstruction performance of the DPS method compared to FBPConvNet may be attributed to their different training objectives. The DPS method utilizes a loss function that aims to minimize the expected Wasserstein-1 distance between the distribution provided by the generator and the posterior distribution of the underlying image given measurements, which typically encourages the generator to output sharper and more realistic textures. On the other hand, FBPConvNet in our experiments was trained with the mean squared error loss, which is known to promote smooth reconstructions and can lead to the loss of fine details. In contrast, in the MRI experiments, we observe that BPConvNet achieves better reconstruction performance than all UQVAE instances. In our experiments, we used only the baseline configuration of the UQVAE method, i.e., without adversarial loss and without recurrent blocks. We believe this choice may explain the observed result, since those two components have been shown to provide improvements in reconstruction performance [66].
Figure 5: SSIM results for filtered backprojection (FBP), FBPConvNet [75] (FBPCN), deep posterior sampling [22] (DPS), zero-filling (ZF), BPConvNet [75] (BPCN), the variational autoencoder-based posterior sampling method [66] (UQVAE), and the proposed deep ensembling [41]-based framework (DE). Results shown for the CT (top) and MRI (bottom) problems.
@@ III-F Computationally Efficient Ensembling
Figure 6: SSIM and negative predictive log-likelihood (NPLL) results for diffusion posterior sampling [14] (DiPS), proposed framework with MC Dropout ensembling (MC-DC and MC-DE), deep ensembling (DE), and SWAG ensembling (SLE). The results are presented for the image inpainting problem described in Section III-A.